eighty liters of a 40% antifreeze solution had to be strengthened so it contained 52% antifreeze. how much should be added?
Answer by Nate(3500)

I'm guessing you are adding 100% or pure antifreeze.
( 80 * 40% + x * 100% ) / ( 80 + x ) = 52%
( 32 + x ) / ( 80 + x ) = 0.52
32 + x = 0.52( 80 + x )
32 + x = 41.6 + 0.52x
0.48x = 9.6
x = 20 Litres of antifreeze